A 240 V two-bar electrical eadiator is rated 1 KW per bar. Calculate the resistance in the radiator circuit when one bar is switched on.


1
Expert's answer
2022-03-13T18:46:16-0400

Solution.

"V=240V;"

"P=1000W;"

"R-?;"

"P=\dfrac{V^2}{R}\implies R=\dfrac{V^2}{P};"

"R=\dfrac{240^2}{1000}=57.6\Omega;"

Answer: "R=57.6\Omega."